# Compositional amortized inference for large-scale hierarchical Bayesian models

### v1

@@ Abstract
Amortized Bayesian inference (ABI) has emerged as a powerful simulation-based approach for estimating complex mechanistic models, offering fast posterior sampling via generative neural networks. However, extending ABI to hierarchical models, a cornerstone of modern Bayesian analysis, remains a major challenge due to the difficulty of scaling to large numbers of parameters. In this work, we build on compositional score matching (CSM), a divide-and-conquer strategy for Bayesian updating using diffusion models. To address existing stability issues of CSM, we propose adaptive solvers coupled with a novel, error-damping compositional estimator. Our proposed method remains stable even with hundreds of thousands of data points and parameters. We validate our approach on a controlled toy example, a high-dimensional spatial autoregressive model, and a real-world advanced microscopy biological application task involving over 750,000 parameters.
@@ 1 Introduction
Simulation-based inference [SBI; Diggle and Gratton, 1984, Cranmer et al., 2020] is entering a new era, leveraging deep learning advances to deliver markedly more efficient computational statistics. Within this framework, amortized Bayesian inference [Bürkner et al., 2023, ABI;] now scales Bayesian analysis to high-dimensional, mechanistic models, driving state-of-the-art discoveries in fields as diverse as astrophysics [Dax et al., 2025] and neuroscience [Tolley et al., 2024].
@@ 2 Background and related work
Hierarchical Bayesian models are the default choice to model dependencies in nested data, where observations are organized into clusters or levels [Gelman et al., 2013]. From a Bayesian perspective, any parametric data model $p(\mathbf{Y}\mid\boldsymbol{\theta})$ can incorporate multilevel structure via a hierarchical prior. For instance, a two-level model defines two generative stages
@@
via a hyperprior $p(\boldsymbol{\eta})$ capturing global variation and a conditional prior $p(\boldsymbol{\theta}\mid\boldsymbol{\eta})$ capturing local variation. The task of Bayesian estimation is to estimate the full joint posterior over local and global parameters:
@@
$\min_{q}\left\{\mathbb{E}_{p(\boldsymbol{\theta},\mathbf{Y})}\Big{[}\mathcal{S}(q(\cdot\mid\mathbf{Y}),\boldsymbol{\theta})\Big{]}\approx\frac{1}{B}\sum_{b=1}^{B}\mathcal{S}(q(\cdot\mid\mathbf{Y}^{(b)}),\boldsymbol{\theta}^{(b)})\right\}.$ (3)
@@
Scalability issues arise since the expectation in Eq. 3 now runs over $p(\boldsymbol{\eta},\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{J},\mathbf{Y}_{1},\dots,\mathbf{Y}_{J})$, necessitating the simulation of a data set of data sets $\{\mathbf{Y}_{1},\dots,\mathbf{Y}_{J}\}$ for each batch instance (cf. Figure 1, left): even for $J\approx 1000$, a single training batch requires tens of thousands of simulations, exceeding typical simulation budgets for non-trivial models quickly. Similar-sized problems can also become practically infeasible for established MCMC samplers [e.g., NUTS Hoffman et al., 2014] even for models with closed-form likelihoods (see Experiment 2).
Inspired by Geffner et al. [2023], Linhart et al. [2024], Gloeckler et al. [2024b], we address these efficiency issues in a “divide-and-conquer” manner via compositional score matching (CSM; cf. Figure 1, right). Along the way, we introduce several key improvements to CSM in terms of stability and scalability. To the best of our knowledge, we provide the first simulation-based method capable of handling large-scale hierarchical Bayesian models with or without explicit likelihoods.
@@
Score-based modeling [Song and Ermon, 2019] and diffusion models [Ho et al., 2020] provide a powerful framework for generative modeling by learning to reverse a noise-adding process. In a non-hierarchical setting, diffusion models define a forward process that gradually corrupts a sample $\boldsymbol{\theta}$ into pure noise, at each time step $t\in[0,1]$ typically taking the form
$\boldsymbol{\theta}_{t}=\alpha_{t}\boldsymbol{\theta}+\sigma_{t}\boldsymbol{\epsilon}\quad\text{with}\quad\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I}).$
The factors $\alpha_{t}$ and $\sigma_{t}$ are time-dependent functions that satisfy $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ for variance-preserving processes. These functions are often parameterized in terms of the log signal-to-noise ratio $\lambda_{t}=\log(\alpha_{t}^{2}/\sigma_{t}^{2})$, known as the noise schedule of the forward process [Kingma and Gao, 2023].
The conditional denoising score matching loss can be expressed in terms of an unconditional score, as shown by Li et al. [2024], and further reformulated as an $\boldsymbol{\epsilon}$-prediction objective following Kingma and Gao [2023]:
@@
which assumes the equivalent score parameterization $\hat{\boldsymbol{\epsilon}}=-s_{\boldsymbol{\psi}}(\boldsymbol{\theta}_{t},\mathbf{Y},\lambda_{t})\sigma_{t}$. The weighting function $w_{t}>0$, often chosen to match the noise schedule $\lambda_{t}$ (see Kingma and Gao [2023] for a detailed review of different weighting functions and noise schedules), is instantiated here as the likelihood weighting proposed by Song et al. [2021]. The forward and backward diffusion process can be specified as a stochastic differential equation [SDE; Song et al., 2020], which enables posterior sampling using state-of-the-art SDE solvers (more details in Appendix A.1). Moreover, it has neither been used for hierarchical modeling nor explored for compositional score matching in prior work Geffner et al. [2023], Linhart et al. [2024], Gloeckler et al. [2024b], as discussed next.
Table 1: Convergence of sampling methods for Gaussian toy example (Experiment 1) with a maximum budget of 10,000 compositional sampling steps (✓– converges, ✗– fails).
Method $N{=}10$ $N{=}100$ $N{=}10\text{k}$ $N{=}100\text{k}$
Annealed Langevin sampler
@@ 3.1 Compositional score matching
A major challenge in Bayesian inference arises when dealing with varying and potentially large numbers of observations, especially in hierarchical models. To address this for non-hierarchical models, Geffner et al. [2023] introduced compositional score matching (CSM) that enables the combination of multiple conditionally independent score estimates into a global posterior approximation.
Suppose we have $J$ exchangeable groups $\{\mathbf{Y}_{j}\}_{j=1}^{J}$, then we can divide the groups in $k=\lfloor{J/m}\rfloor$ subsets of size $m$, where we denote the group in subset $i$ with an additional index $l_{i}$. Then, the compositional posterior can be written as $p(\boldsymbol{\theta}\mid\{\mathbf{Y}_{j}\}_{j=1}^{J})=p(\boldsymbol{\theta})^{1-k}\prod_{i=1}^{k}p(\boldsymbol{\theta}\mid\{\mathbf{Y}_{l_{i}}\}_{l_{i}=1}^{m})$ using Bayes’ rule twice. Here, $m$ determines the number of simulations we need to perform to generate one sample for training as we condition the score on multiple groups. The larger $m$ gets, the more simulations for training are needed, but this might reduce approximation error of the score. As a bridging density between the latent diffusion prior, a spherical Gaussian $p_{t=1}(\boldsymbol{\theta})=\mathcal{N}\left(\mathbf{0},\frac{1}{k}\mathbf{I}\right)$, and the posterior $p_{t=0}(\boldsymbol{\theta}\mid\{\mathbf{Y}_{j}\}_{j=1}^{J})$, Geffner et al. [2023] choose a linear interpolation between the prior and the individual posterior scores
$\nabla_{\boldsymbol{\theta}}\log p_{t}(\boldsymbol{\theta}\mid\{\mathbf{Y}_{j}\}_{j=1}^{J})=(1-k)(1-t)\nabla_{\boldsymbol{\theta}}\log p(\boldsymbol{\theta})+\sum_{i=1}^{k}\nabla_{\boldsymbol{\theta}}\log p_{t}(\boldsymbol{\theta}\mid\{\mathbf{Y}_{l_{i}}\}_{l_{i}=1}^{m}).$ (5)
This allows us to train a single score network on samples $(\boldsymbol{\theta}_{i},\{\mathbf{Y}_{l_{i}}\}_{l_{i}=1}^{m})\sim p(\boldsymbol{\theta})p(\{\mathbf{Y}_{l_{i}}\}_{l_{i}=1}^{m}\mid\boldsymbol{\theta})$. After training, we can perform inference by sampling from the latent prior $p_{t=1}(\boldsymbol{\theta})=\mathcal{N}\left(\mathbf{0},\frac{1}{k}\mathbf{I}\right)$ and then using the compositional score to sample from the posterior.
@@
Geffner et al. [2023] employ a Langevin sampling scheme to invert the diffusion process for posterior inference, which needs many score evaluations for accurate inference [Jolicoeur-Martineau et al., 2021] and is sensitive to the choices of step-size at each sampling iteration [Linhart et al., 2024]. In contrast, Linhart et al. [2024] propose a second order approximation of the backward diffusion kernels to bypass the need for Langevin dynamics, however introducing the need to approximate potentially large covariance matrices, limiting their experiments to only 100 observations.
@@
In the remainder, we demonstrate that it is possible to leverage the SDE formulation by using the compositional score in the reverse SDE (see Appendix A.1) to sample from the posterior distribution. This allows us to use more efficient numerical solvers. However, regardless of the number of conditioning groups $k$, increasing the number of score terms leads to error compounding, resulting in unstable dynamics and divergent samples (see Figure 6 in Appendix). Even higher-order solvers require extremely small step sizes, constraining their practicality (cf. Table 1). To address this, we introduce an error-damping, unbiased mini-batch estimator and demonstrate its scalability.
3.2 Compositional hierarchical score matching
In this study, we propose to approximate the posterior distributions at various levels of a hierarchical model using score matching. Each level is associated with its own neural estimator, responsible for modeling the corresponding posterior distribution. This is similar to the framework introduced by Habermann et al. [2024] and Heinrich et al. [2023] but avoids the need for hierarchical embeddings.
At each level, the posterior is modeled via a conditional score function. At higher levels of the hierarchy, we leverage the compositional score framework, enabling training of the global model on single groups. For example, in a two level model, we have
$s_{\boldsymbol{\psi}_{\text{local}}}(\boldsymbol{\theta}_{j},\boldsymbol{\eta},\mathbf{Y}_{j},\lambda_{t})\approx\nabla_{\boldsymbol{\theta}_{j}}\log p_{t}(\boldsymbol{\theta}_{j}\mid\boldsymbol{\eta},\mathbf{Y}_{j}),\quad s_{\boldsymbol{\psi}_{\text{global}}}(\boldsymbol{\eta},\mathbf{Y}_{j},\lambda_{t})\approx\nabla_{\boldsymbol{\eta}}\log p_{t}(\boldsymbol{\eta}\mid\mathbf{Y}_{j}).$ (6)
For each group, we may learn a shared summary representation $\mathbf{h}_{j}=h(\mathbf{Y}_{j})$ via a summary network $h$. Either the raw data $\mathbf{Y}_{j}$ or its summary $\mathbf{h}_{j}$ is then used as input to both the global and local score-based models. The design of the summary $h$ should be adapted to the specific data modality (e.g., recurrent networks or transformers for time series, etc). When conditioning on multiple groups, we utilize exchangeability via a second summary network (e.g., a Deep Set model [Zaheer et al., 2017]), which aggregates the individual summaries into a permutation-invariant global summary.
Both the global and local score networks can be trained jointly via denoising score matching objectives, allowing the system to share representations and learn in a coherent hierarchical fashion:
$\min_{\boldsymbol{\psi}}\mathbb{E}_{p(\boldsymbol{\theta},\boldsymbol{\eta},\mathbf{Y})}\mathbb{E}_{t\sim\mathcal{U}(0,1)}w_{t}\left[\|\boldsymbol{\epsilon}+s_{\boldsymbol{\psi}_{\text{local}}}(\boldsymbol{\theta}_{t},\boldsymbol{\eta},\mathbf{Y},\lambda_{t})\sigma_{t}\|_{2}^{2}+\|\boldsymbol{\epsilon}+s_{\boldsymbol{\psi}_{\text{global}}}(\boldsymbol{\eta}_{t},\mathbf{Y},\lambda_{t})\sigma_{t}\|_{2}^{2}\,\right],$ (7)
with $\boldsymbol{\eta}_{t}=\alpha_{t}\boldsymbol{\eta}+\sigma_{t}\boldsymbol{\epsilon}$ and $\boldsymbol{\theta}_{t}=\alpha_{t}\boldsymbol{\theta}+\sigma_{t}\boldsymbol{\epsilon}$, where $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$. Having trained a global and a local model, we can sample from the joint posterior by
$\boldsymbol{\eta}^{(s)}$ $\sim q_{\boldsymbol{\psi}_{\text{global}}}(\boldsymbol{\eta}\mid\{\mathbf{Y}_{j}\}_{j=1}^{J}),\qquad\boldsymbol{\theta}^{(s)}_{j}\sim q_{\boldsymbol{\psi}_{\text{local}}}(\boldsymbol{\theta}\mid\boldsymbol{\eta}^{(s)},\mathbf{Y}_{j}),$ (8)
where we leverage the compositional score for the global parameters and then sample the local ones.
3.3 Scalable compositional score matching
@@
We propose using the SDE formulation to perform compositional inference with adaptive solvers—avoiding the need for annealed Langevin sampling, which requires many steps per noise level and becomes prohibitively expensive when error correction is needed. This allows us to leverage an adaptive solver that automatically adjusts the step size during integration, which is especially beneficial for large values of $J$, where the need for finer granularity (i.e., smaller step sizes) increases and manual tuning becomes infeasible [Jolicoeur-Martineau et al., 2021].
However, two main challenges remain in scaling the compositional approach to large datasets: 1) the bridging densities introduced by Geffner et al. [2023] become unstable as $J$ increases (see Table 1) and 2) the memory requirements grow substantially when accumulating scores over the full data set.
Flexible error-dampening bridging densities
@@
The stability of the bridge density might be improved by introducing a damping factor of the accumulated score. Yet, naively applying a dampening factor to the compositional score to prevent it from diverging would bias the posterior sample. Instead, to mitigate instability at large $J$, we introduce a more flexible class of error-dampening bridging densities:
$p_{t}(\boldsymbol{\theta}\mid\{\mathbf{Y}_{j}\}_{j=1}^{J})\propto\left(p(\boldsymbol{\theta})^{1-J}\right)^{(1-t)d(t)}\prod_{j=1}^{J}p_{t}(\boldsymbol{\theta}\mid\mathbf{Y}_{j})^{d(t)},$ (9)
@@
where $d(0)=d_{0}=1$ and $d(1)=d_{1}\leq 1$ and latent diffusion prior $p_{t=1}(\boldsymbol{\theta})=\mathcal{N}(0,\frac{1}{nd_{1}}\mathbf{I})$.
@@
The key idea is to define a monotonic function $d(t)$ that moderates the accumulation of score contributions throughout the diffusion trajectory. In high-noise regimes, we reduce the influence of the individual terms to prevent the score from diverging, while for $t\to 0$, we allow their contributions to accumulate, recovering the true posterior. This construction is motivated by the observation that adaptive solvers require smaller steps in high-noise regimes to avoid numerical instability (see Appendix Figure 6). As damping schedule, we use an exponential decay $d(t)=d_{0}\cdot\exp(-\ln(d_{0}/d_{1})\cdot t)$ with $d_{0}=1$ and a hyperparameter $d_{1}$ that can be tuned during inference.
@@
$\hat{s}_{\boldsymbol{\psi}}(\boldsymbol{\theta},\{\mathbf{Y}_{j}\}_{j=1}^{J},\lambda_{t})=(1-J)(1-t)\nabla_{\boldsymbol{\theta}}\log p(\boldsymbol{\theta})+\frac{J}{M}\sum_{i=1}^{M}s_{\boldsymbol{\psi}}(\boldsymbol{\theta},\mathbf{Y}_{j_{i}},\lambda_{t}),$ (10)
@@
The mini-batch estimator in (10) is an unbiased estimate of the compositional score.
For a short proof, see Appendix A.2. Combining this estimator with the dampening function yields our final form:
$\hat{s}^{d}_{\boldsymbol{\psi}}(\boldsymbol{\theta},\{\mathbf{Y}_{j}\}_{j=1}^{J},\lambda_{t})=d(t)\cdot\left((1-J)(1-t)\nabla_{\boldsymbol{\theta}}\log p(\boldsymbol{\theta})+\frac{J}{M}\sum_{i=1}^{M}s_{\boldsymbol{\psi}}(\boldsymbol{\theta},\mathbf{Y}_{j_{i}},\lambda_{t})\right).$ (11)
This error-damping mini-batch estimator scales well with dataset size and maintains stability across the reverse-time diffusion process.

### v3

@@ Abstract
Amortized Bayesian inference (ABI) has emerged as a powerful simulation-based approach for estimating complex mechanistic models, offering fast posterior sampling via generative neural networks. However, extending ABI to hierarchical models, a cornerstone of modern Bayesian analysis, remains a major challenge due to the need to simulate massive data sets and estimate thousands of parameters. In this work, we build on compositional score matching (CSM), a divide-and-conquer strategy for Bayesian updating using diffusion models. To address existing stability issues of CSM in dealing with large data sets, we couple adaptive solvers with a novel, error-damping compositional estimator. Our estimator remains stable even with hundreds of thousands of data points and parameters. We validate our approach on a controlled toy example, a high-dimensional autoregressive model, and a real-world advanced microscopy application involving over 750,000 parameters.
@@ 1 Introduction
Simulation-based inference [SBI; Diggle and Gratton, 1984, Cranmer et al., 2020] is entering a new era, leveraging deep learning advances to deliver markedly more efficient computational statistics. Within this framework, amortized Bayesian inference [ABI; Bürkner et al., 2023] now scales Bayesian analysis to high-dimensional, mechanistic models, driving state-of-the-art discoveries in fields as diverse as astrophysics [Dax et al., 2025] and neuroscience [Tolley et al., 2024].
@@ 2 Background and related work
Hierarchical Bayesian models are the default choice to model dependencies in nested data, where observations are organized into clusters, levels, or groups [Gelman et al., 2013]. From a Bayesian perspective, any parametric data model $p(\mathbf{Y}\mid\boldsymbol{\theta})$ can incorporate multilevel structure via a hierarchical prior. For instance, a two-level model defines two stages
@@
via a hyperprior $p(\boldsymbol{\eta})$ encoding global variation between groups and a conditional prior $p(\boldsymbol{\theta}\mid\boldsymbol{\eta})$ encoding local variation within groups. The task of Bayesian estimation is to estimate the full joint posterior over local and global parameters:
@@
$\min_{q}\bigg\{\mathbb{E}_{p(\boldsymbol{\theta},\mathbf{Y})}\Big[\mathcal{S}(q(\cdot\mid\mathbf{Y}),\boldsymbol{\theta})\Big]\approx\frac{1}{B}\sum_{b=1}^{B}\mathcal{S}(q(\cdot\mid\mathbf{Y}^{(b)}),\boldsymbol{\theta}^{(b)})\bigg\}.$ (3)
@@
Scalability issues arise since the expectation in Eq. 3 now runs over $p(\boldsymbol{\eta},\boldsymbol{\theta}_{1},\dots,\boldsymbol{\theta}_{J},\mathbf{Y}_{1},\dots,\mathbf{Y}_{J})$, necessitating the simulation of a data set of data sets $\{\mathbf{Y}_{1},\dots,\mathbf{Y}_{J}\}$ for each batch instance (cf. Figure 1, left): even for $J\approx 1000$, a single training batch requires tens of thousands of simulations, exceeding typical simulation budgets for non-trivial models quickly. Similar-sized problems can also become practically infeasible for established MCMC samplers [e.g., NUTS, Hoffman et al., 2014], even for models with closed-form likelihoods (see Experiment 2).
Building on prior work by Geffner et al. [2023], Linhart et al. [2024], Gloeckler et al. [2024b], we address these efficiency issues in a “divide-and-conquer” manner via compositional score matching (CSM; cf. Figure 1, right). Along the way, we introduce several key improvements to CSM in terms of stability and scalability. To the best of our knowledge, we provide the first simulation-based method capable of handling large-scale hierarchical Bayesian models with or without explicit likelihoods.
@@
Score-based modeling [Song and Ermon, 2019] and diffusion models [Ho et al., 2020] provide a powerful framework for generative modeling by learning to reverse a noise-adding process. Diffusion models build on a forward process that gradually corrupts a sample $\boldsymbol{\theta}$ into pure noise at each time step $t\in[0,1]$, typically taking the form $\boldsymbol{\theta}_{t}=\alpha_{t}\boldsymbol{\theta}+\sigma_{t}\boldsymbol{\epsilon}$ with $\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I})$. The factors $\alpha_{t}$ and $\sigma_{t}$ are time-dependent functions that satisfy $\alpha_{t}^{2}+\sigma_{t}^{2}=1$ for variance-preserving processes. These functions are often parameterized in terms of the log signal-to-noise ratio $\lambda_{t}=\log(\alpha_{t}^{2}/\sigma_{t}^{2})$, known as the noise schedule [Kingma and Gao, 2023].
The conditional denoising score matching loss can be expressed in terms of an unconditional score [Li et al., 2024], and further reformulated as an $\boldsymbol{\epsilon}$-prediction objective [Kingma and Gao, 2023]:
@@
which assumes the equivalent score parameterization $\hat{\boldsymbol{\epsilon}}=-s_{\boldsymbol{\psi}}(\boldsymbol{\theta}_{t},\mathbf{Y},\lambda_{t})\sigma_{t}$. The weighting function $w_{t}>0$, often chosen to match the noise schedule $\lambda_{t}$ (see Kingma and Gao [2023] for a detailed review of different weighting functions and noise schedules), is instantiated here as the likelihood weighting proposed by Song et al. [2021]. The forward and backward diffusion process can be specified as a stochastic differential equation [SDE; Song et al., 2020], which enables posterior sampling using state-of-the-art SDE solvers (more details in Appendix A.1). Moreover, this formulation has neither been used for hierarchical modeling nor explored for compositional score matching in most prior work [Geffner et al., 2023, Linhart et al., 2024], as discussed next.
Table 1: Convergence of sampling methods for Gaussian toy example (Experiment 1) with a maximum budget of 10,000 compositional sampling steps and maximal $30\text{\,}\min$ of runtime for a single dataset (✓– converges, ✗– fails).
Method N=10N{=}10 N=100N{=}100 $N{=}10\text{k}$ $N{=}100\text{k}$
Annealed Langevin sampler [Geffner et al., 2023]
@@ 3.1 Compositional score matching
A major challenge in Bayesian inference arises when dealing with varying and potentially large numbers of observations, especially in hierarchical models. To address this for non-hierarchical models, Geffner et al. [2023] introduced compositional score matching (CSM) that enables the aggregation of multiple conditionally independent score estimates into a global posterior estimate. In the following, we first introduce a naive extension of CSM for estimating the global parameters $\boldsymbol{\eta}$ of hierarchical models. We then propose a solution to the stability problems of the naive approach that allows us to estimate the full joint posterior (Eq. 1) of large hierarchical models.
Compositional score and bridging density
Suppose we have $J$ exchangeable groups of data points, $\smash{\{\mathbf{Y}_{j}\}_{j=1}^{J}}$. Then, the compositional posterior can be written as
$p(\boldsymbol{\eta}\mid\{\mathbf{Y}_{j}\}_{j=1}^{J})\propto p(\boldsymbol{\eta})^{1-J}\prod_{j=1}^{J}p(\boldsymbol{\eta}\mid\mathbf{Y}_{j}),$ (5)
using Bayes’ rule twice. Let $p_{t}(\boldsymbol{\eta}_{t}\mid\mathbf{Y}_{j})$ be the time-varying density of the noise-corrupted parameter $\boldsymbol{\eta}_{t}$ for $t\in[0,1]$, as defined by the forward diffusion process. Then, we can define the bridging densities $p_{t}(\boldsymbol{\eta}_{t}\mid\{\mathbf{Y}_{j}\}_{j=1}^{J})\propto p(\boldsymbol{\eta}_{t})^{(1-J)(1-t)}\prod_{j=1}^{J}p_{t}(\boldsymbol{\eta}_{t}\mid\mathbf{Y}_{j})$. This results in a linear composition of the prior score and individual posterior scores:
$\nabla_{\boldsymbol{\eta}_{t}}\log p_{t}(\boldsymbol{\eta}_{t}\mid\{\mathbf{Y}_{j}\}_{j=1}^{J})=(1-J)(1-t)\nabla_{\boldsymbol{\eta}_{t}}\log p(\boldsymbol{\eta}_{t})+\sum_{j=1}^{J}\nabla_{\boldsymbol{\eta}_{t}}\log p_{t}(\boldsymbol{\eta}_{t}\mid\mathbf{Y}_{j}).$ (6)
After training, we can sample from the base distribution $p_{t=1}(\boldsymbol{\eta}_{t})=\mathcal{N}\left(\mathbf{0},\frac{1}{J}\mathbf{I}\right)$ and use the compositional score to sample from the posterior distribution of $\boldsymbol{\eta}$. The score model can also be conditioned on $m$ groups jointly, rather than a single group. This results in a compositional update that involves only $k=\lfloor{J/m}\rfloor$ scores per posterior evaluation, which can improve the robustness of the score estimation. However, this comes at an increased computational cost because each training iteration requires a batch simulation of $m$ full groups of data points, rather than just one group.
Sampling with compositional scores
@@
Geffner et al. [2023] employ annealed Langevin sampling to invert the diffusion process for posterior inference, which needs many score evaluations for accurate inference [Jolicoeur-Martineau et al., 2021] and is sensitive to the choice of step-size at each sampling iteration. In contrast, Linhart et al. [2024] use a second-order Gaussian approximation of the backward diffusion kernels to bypass Langevin sampling, introducing the need to approximate potentially large covariance matrices and limiting their experiments to only 100 observations.
@@
In the remainder, we demonstrate that it is possible to instead leverage the SDE formulation by aggregating the compositional scores in the reverse SDE (see Appendix A.1) to sample from the posterior. This allows us to use more efficient numerical solvers. However, regardless of the number of conditioning groups $k$, increasing the number of score terms leads to error compounding due to a potential mismatch of marginal densities $p_{t}$ and the corresponding forward diffusion process, resulting in unstable dynamics and divergent samples (see Figure 5 in Appendix). Even higher-order solvers require extremely small step sizes, constraining their practicality (cf. Table 1). The next section introduces three new modifications to the naive CSM approach that stabilize the bridging density (Eq. 6) and unlock unprecedented scalability to large data sets.
3.2 Stabilizing and scaling compositional score matching
@@
In contrast to most previous work, we adopt the SDE formulation to perform compositional inference with adaptive solvers that automatically adjusts the step size during integration. This modification is essential for larger numbers of groups $J$, where the need for finer granularity (i.e., smaller step sizes) increases and manual tuning becomes infeasible [Jolicoeur-Martineau et al., 2021]. Moreover, it avoids the need for annealed Langevin sampling, which requires many steps per noise level and becomes prohibitively expensive when error correction is needed. However, simply using an adaptive solver does not address the two major challenges for scaling the compositional approach to very large datasets: 1) the bridging densities (Eq. 6) become unstable as $J$ increases (see Table 1) and 2) the memory requirements grow substantially when accumulating scores over the full data set.
Flexible error-damping bridging densities
@@
We propose to stabilize the bridging density by introducing a damping factor of the accumulated score. Yet, naively applying a damping factor to the compositional score to prevent it from diverging would bias the posterior samples. Instead, to mitigate instability at large $J$, we introduce a more flexible class of error-damping bridging densities:
$p_{t}(\boldsymbol{\eta}_{t}\mid\{\mathbf{Y}_{j}\}_{j=1}^{J})\propto\left(p(\boldsymbol{\eta}_{t})^{1-J}\right)^{(1-t)d(t)}\prod_{j=1}^{J}p_{t}(\boldsymbol{\eta}_{t}\mid\mathbf{Y}_{j})^{d(t)},$ (7)
@@
where $d(0)=d_{0}=1$ and $d(1)=d_{1}\leq 1$, and the latent diffusion prior is $p_{t=1}(\boldsymbol{\eta}_{1})=\mathcal{N}(\mathbf{0},\frac{1}{Jd_{1}}\mathbf{I})$.
@@
The key idea is to define a monotonic function $d(t)$ that modulates the accumulation of score contributions throughout the diffusion trajectory during inference. In high-noise regimes, we reduce the influence of the individual terms to prevent the score from diverging, while for $t\to 0$, we allow their contributions to accumulate, recovering the true posterior. This construction is motivated by the observation that adaptive solvers require smaller steps in high-noise regimes to avoid numerical instability (see Appendix Figure 5). As a damping schedule, we use an exponential decay $d(t)=d_{0}\cdot\exp(-\ln(d_{0}/d_{1})\cdot t)$ with $d_{0}=1$ and a hyperparameter $d_{1}$ that can be tuned during inference.
@@
$\hat{s}_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\{\mathbf{Y}_{j}\}_{j=1}^{J},\lambda_{t})=(1-J)(1-t)\nabla_{\boldsymbol{\eta}_{t}}\log p(\boldsymbol{\eta}_{t})+\frac{J}{M}\sum_{i=1}^{M}s_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\mathbf{Y}_{j_{i}},\lambda_{t}),$ (8)
@@
The mini-batch estimator in (8) is an unbiased estimator of the compositional score.
For a short proof, see Appendix A.2. Combining this estimator with the damping function yields our final compositional form:
$\hat{s}^{d}_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\{\mathbf{Y}_{j}\}_{j=1}^{J},\lambda_{t})=d(t)\cdot\bigg((1-J)(1-t)\nabla_{\boldsymbol{\eta}_{t}}\log p(\boldsymbol{\eta}_{t})+\frac{J}{M}\sum_{i=1}^{M}s_{\boldsymbol{\psi}}(\boldsymbol{\eta}_{t},\mathbf{Y}_{j_{i}},\lambda_{t})\bigg).$ (9)
This error-damping mini-batch estimator scales well with increasing numbers of groups $J$ and maintains stability across the reverse-time diffusion process, as shown in our experiments.